# Generalized hydrodynamics of integrable quantum circuits

### v1

@@ I Introduction
Figure 1: Sketch of the bipartition protocol. The system is initialized by joining at the origin two product states (the figure corresponds to the Néel and anti-Néel states) and evolved by applying a homogeneous quantum circuit, Trotterizing the integrable XXZ continuous-time dynamics ($\tau$ is the duration of the Trotter step). A lightcone spreads ballistically from the junction, inside of which non-trivial dynamics takes place. We study the profiles of local observables arising at late times. The plot shows the example of the staggered spin-$z$ magnetization and is obtained by solving the GHD equations, cf. Sec. IV.
Integrable quantum circuits providing a Trotterization of integrable Hamiltonians were introduced in Refs. [29, 30, 31]. In the past few years they have received significant attention due to their dynamical features [32, 33, 34], their rich underlying mathematical structures [35, 36, 37, 38] and seminal work demonstrating experimental implementation in current quantum devices [39, 40, 41]. Prominently, Ref. [33] has shown the existence of Trotter transitions in integrable quantum circuits, manifesting in the correlations of the so-called generalized Gibbs ensemble (GGE) [26, 27, 28]. As GGEs describe the late-time physics of integrable systems, this result is interesting, as it implies that the information on the discretized structure of the dynamics is not washed out at coarse-grained time scales.
@@
The rest of this work is organized as follows. We begin in Sec. II introducing the XXZ quantum circuit and its Bethe ansatz solution. In Sec. III we discuss the bipartition protocol and derive the necessary technical ingredients for its analytic description. They include obtaining the GGEs corresponding to different initial states, generalizing the results of Ref. [33], and deriving the GHD equations for integrable quantum circuits. Our results are discussed in detail in Sec. IV, where we provide predictions for the system dynamics at different points in the parameter space, by numerically solving the GHD equations. Our conclusions are presented in Sec. V, while the appendices contain the most technical aspects of our work.
II The model
II.1 The integrable quantum circuit
@@
$𝟙V_{n,n+1}(\tau)=e^{-i\frac{\tau}{4}\left[\sigma_{n}^{x}\sigma_{n+1}^{x}+\sigma_{n}^{y}\sigma_{n+1}^{y}+\Delta(\sigma_{n}^{z}\sigma_{n+1}^{z}-\openone)\right]}\,,$ (4)
@@
$𝟙H=\frac{1}{4}\sum_{j=1}^{L}\left[\sigma^{x}_{j}\sigma^{x}_{j+1}+\sigma^{y}_{j}\sigma^{y}_{j+1}+\Delta(\sigma^{z}_{j}\sigma^{z}_{j+1}-\openone)\right]\,,$ (5)
@@
II.2 The Bethe ansatz solution
@@
[…]
@@ II.2.1 The TBA in the gapped regime
$a_{nm}(\lambda)=$ $(1-\delta_{nm})a_{|n-m|}(\lambda)+2a_{|n-m|+2}(\lambda)$
$+$ $\ldots+2a_{n+m-2}(\lambda)+a_{n+m}(\lambda)\,,$ (14)
@@
$J_{2j}$ $={\mathcal{N}_{x}}^{-1}\left[2(\sigma_{2j-1}^{+}\sigma_{2j}^{-}-\sigma_{2j-1}^{-}\sigma_{2j}^{+})\right.$
$+$ $\left.i\sinh(ix)(\sigma_{2j-1}^{z}-\sigma_{2j}^{z})\right]\,,$ (22)
@@ II.2.2 The TBA in the gapless regime
$\mathrm{sign}(q_{m})\left[\rho_{m}^{t}(\lambda)\right]=$ $a_{m}^{(ix/2)}(\lambda)$
-- $\sum_{n=1}^{N_{b}}\left(a_{mn}\ast\rho_{m}\right)(\lambda)\,,$ (32)
@@
$a_{jk}(\lambda)$ $=(1-\delta_{n_{j}n_{k}})a_{|n_{j}-n_{k}|}^{\upsilon_{j}\upsilon_{k}}(\lambda)+2a_{|n_{j}-n_{k}|+2}^{\upsilon_{j}\upsilon_{k}}(\lambda)+\ldots$
$+$ $2a_{n_{j}+n_{k}-2}^{\upsilon_{j}\upsilon_{k}}(\lambda)+a_{n_{j}+n_{k}}^{\upsilon_{j}\upsilon_{k}}(\lambda)\,.$ (35)
@@
$\mathrm{sign}(q_{n})\left[\rho_{n}^{t}(\lambda)v^{\rm eff}_{n}(\lambda)\right]=$ $\tfrac{1}{2\pi}\mathrm{sign}(q_{n})E^{\prime}_{n}(\lambda)$
-- $\sum_{m=1}^{N_{b}}\left(a_{nm}\ast(\rho_{m}v^{\rm eff}_{m})\right)(\lambda)\,,$ (36)
@@
$a^{(m)\,{\rm dr}}_{j}(\lambda)$ $=a_{p}(\lambda-(-1)^{m}\tfrac{ix}{2})$
-- $\sum_{q=1}^{N_{s}}\int{{\rm d}\mu}\,a_{jq}(\lambda-\mu)\theta_{q}(\mu)\sigma_{q}a^{(m)\,{\rm dr}}_{q}(\mu),$ (40)
@@
III The bipartition protocol
III.1 The initial states
@@
Alternatively, one could consider preparing two halves of the system in non-equilibrium states with a simple structure [48]. As discussed in Sec. I, each half of the system far enough from the junction is then expected to relax to a stationary GGE which depends on the initial state. Accordingly, at late time the non-trivial dynamics again takes place only in a region increasing linearly in time and centered around the junction.
@@ III.1.1 Gapped phase
III.1.2 Gapless phase
As stated previously, the GGE description in the gapless phase depends in a non-trivial way on the specific value of $\gamma$. Following Section II.2, we shall limit ourselves to the simplest case where $\gamma$ takes the form (28). The GGE is then characterized by a family of functions $\{\eta_{j}\}_{j=1,\ldots p+1}$, which are obtained from a set of auxiliary functions $\{Y_{j}\}_{j=1,\ldots p}$, $K$, $\mathfrak{b}$ whose construction is detailed in Appendix B. For concreteness, we give below the explicit rapidity distributions for $p=2,3$. For $p=2$ and the Néel state, we have
@@ III.2 GHD equations
As we will discuss in the next subsection, the form of the scattering phase depends on the phase of the model and it can be read off by comparing  (III.2) with Eqs. (18) and (36), respectively.
III.3 The quasi-momentum and the quasi-energy
@@
III.3.1 Gapped phase
As discussed in Sec. II, the eigenvalues of $U(\tau)$ depend on the solutions $\{p_{j}\}$ to the Bethe equation (8). Denoting by $u(\{p_{j}\})$ the eigenvalue of $U(\tau)$ corresponding to a set $\{p_{j}\}_{j=1}^{M}$, we have
@@ III.3.1 Gapped phase
$E_{n}(\lambda)$ $=\!\sum_{j=-\tfrac{n-1}{2}}^{\tfrac{n-1}{2}}\varepsilon(i(\lambda+j\eta))$
$=$ $-\frac{i}{\tau}\!\log\quantity[\frac{\sin(\lambda-\tfrac{x}{2}+in\tfrac{\eta}{2})\sin(\lambda+\tfrac{x}{2}-in\tfrac{\eta}{2})}{\sin(\lambda-\tfrac{x}{2}-in\tfrac{\eta}{2})\sin(\lambda\!+\!\tfrac{x}{2}+in\tfrac{\eta}{2})}].$ (61)
@@
$E^{\prime}_{n}(\lambda)\!$ $=\!\frac{-\sinh({n\eta})/\tau}{\cosh^{2}(\tfrac{n\eta}{2})\sin^{2}(\lambda-\tfrac{x}{2})+\sinh^{2}(\tfrac{n\eta}{2})\cos^{2}(\lambda-\tfrac{x}{2})}$
$\quad-(x\to-x)\,.$ (62)
@@ IV Results
where $\rho^{L}_{n}(\lambda)$, $\rho^{R}_{n}(\lambda)$ are the GGE distribution functions associated with the left and right initial states. For the product states considered in this work, these functions were computed in Sec. III. The above boundary conditions encode the fact that, away from the junction point, the system will locally equilibrate to a homogeneous GGE, determined by the initial states.
Solving (71) with the boundary conditions (72) can be done with standard iterative methods, see e.g. Ref. [94]. In this work, we use a different numerical approach which was recently introduced in Ref. [95] and briefly explained in Appendix A. In the rest of this section we present the final results of our numerical computations. Note that the solution to (71) gives us a family of rapidity distribution functions $\{\rho_{n}(\zeta,\lambda)\}_{\zeta}$. We use the latter to compute the space-time profiles of local observables, using the formulas presented in Sec. III.
IV.1 The non-interacting points
@@
IV.2 Observable profiles at the interacting points
@@ V Outlook
Acknowledgments
We acknowledge helpful discussions with Bruno Bertini and especially Benjamin Doyon for early collaboration on this project. FH acknowledges funding from the faculty of Natural, Mathematical & Engineering Sciences at King’s College London. The CREATE cluster [112] was used for numerical simulations. This work was co-funded by the European Union (ERC, QUANTHEM, 101114881). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.
@@ Appendix A Numerical solution of the GHD equation
$K_{0;i}(x,\lambda)$ $={P^{\prime}}^{\mathrm{dr}}_{\operatorname{sgn}(x),i}(\lambda)x$ (76)
$\hat{\theta}_{0;i}(k,p)$ $=\theta_{\operatorname{sgn}(k);i}(\lambda)$ (77)
$\hat{N}_{0;i}(k,p)$ $=\theta_{\operatorname{sgn}(k);i}(\lambda)k$ (78)
@@
$=P^{\prime}_{i}(\lambda)x-E_{i}^{\prime}(\lambda)t+\hat{\mathbf{T}}[\theta_{\operatorname{sgn}(K_{j}(t,x,\mu));j}(\mu)K_{j}(t,x,\mu)].$ (79)
@@
$\theta_{i}(t,x,\lambda)=\hat{\theta}_{i}(K_{i}(t,x,\lambda),\lambda)=\theta_{\operatorname{sgn}(K_{i}(t,x,\lambda));i}(\lambda).$ (80)
We solve Eq. (79) by iteration with initial guess $P^{\prime}_{i}(\lambda)x-E^{\prime}_{i}(\lambda)t$, which converges in the gappless phase, where only finitely many strings are present.
@@ A.1 Upgraded algorithm
As it turns out for the Néel and Dimer state in the gapped phase this iterative scheme does not converge. The same holds for the dressing equation (which can be viewed as a linearized version of (79)). The iteration scheme does not converge because $n_{i}(\lambda)$ does not decay sufficiently fast for large strings $i\to\infty$: Consider the first term of the iteration procedure to compute $f^{\mathrm{dr}}_{i}(\lambda)$
$f^{\mathrm{dr}}_{i}(\lambda)$ $=f_{i}(\lambda)+\hat{\mathbf{T}}\theta_{j}(\mu)f_{j}(\mu)+\ldots$ (81)
@@
$f_{i}^{\mathrm{dr}}(\lambda)$ $=f_{i}(\lambda)+\hat{\mathbf{T}}\theta_{j}(\mu)f_{j}^{\mathrm{dr}}(\mu).$ (82)
We can rewrite this as follows: Choose any functions $g_{i}(\lambda)$ and $m_{i}(\lambda)$ and define $b_{i}(\lambda)=g_{i}(\lambda)-\hat{\mathbf{T}}m_{j}(\mu)g_{j}(\mu)$. Now write $f^{\mathrm{dr}}_{i}(\lambda)=w_{i}(\lambda)g_{i}(\lambda)$ and observe:
$w_{i}(\lambda)g_{i}(\lambda)$ $=f_{i}(\lambda)+\hat{\mathbf{T}}\theta_{j}(\mu)w_{j}(\lambda)g_{j}(\lambda)$ (83)
$w_{i}(\lambda)g_{i}(\lambda)$ $=w_{i}(\lambda)b_{i}(\lambda)+w_{i}(\lambda)\hat{\mathbf{T}}m_{j}(\mu)g_{j}(\mu).$ (84)
@@
$w_{i}(\lambda)$ $=\frac{1}{b_{i}(\lambda)}\left(f_{i}(\lambda)+\hat{\mathbf{T}}\theta_{j}(\mu)g_{j}(\lambda)w_{j}(\lambda)\right.$
-- $\left.w_{i}(\lambda)\hat{\mathbf{T}}m_{j}(\mu)g_{j}(\mu)\right).$ (85)
This is a new fixed point equation for $w_{i}(\lambda)$. It is motivated by the equation for the effective velocity, which is obtained in the special case $f_{i}(\lambda)=E^{\prime}_{i}(\lambda),b_{i}(\lambda)={P^{\prime}}_{i}(\lambda)$ and $m_{i}(\lambda)=n_{i}(\lambda)$. Intuitively speaking we need that $\zeta_{i}(\lambda)$ decays fast for large $i$, but at the same time $b_{i}(\lambda)$ should decay slowly, such that (85) converges.
For the gapped phase we used $g_{i}(\lambda)=1/2^{i}$ and $m_{i}(\lambda)=1$, which happened to make (85) convergent with initial guess $f_{i}(\lambda)/b_{i}(\lambda)$.
@@
$\tilde{K}_{i}(t,x,\lambda)$ $=\tfrac{1}{b_{i}(\lambda)}\Big{(}P^{\prime}_{i}(\lambda)x-E_{i}^{\prime}(\lambda)t-\tilde{K}_{i}(t,x,\lambda)\hat{\mathbf{T}}m_{j}(\mu)g_{j}(\mu)+\hat{\mathbf{T}}[\hat{N}_{0;j}(g_{j}(\mu)\tilde{K}_{j}(t,x,\mu),\mu)]\Big{)},$ (86)
@@
$\tilde{K}_{i}(t,x,\lambda)$ $=\tfrac{1}{b_{i}(\lambda)}\Big{(}P^{\prime}_{i}(\lambda)x-E_{i}^{\prime}(\lambda)t-\tilde{K}_{i}(t,x,\lambda)\hat{\mathbf{T}}m_{j}(\mu)g_{j}(\mu)+\hat{\mathbf{T}}[\theta_{\operatorname{sgn}(\tilde{K}_{j}(t,x,\mu));j}(\mu)g_{j}(\mu)\tilde{K}_{j}(t,x,\mu)]\Big{)}.$ (87)

### v2

@@ 1 Introduction
Figure 1: Sketch of the bipartition protocol. The system is initialized by joining at the origin two product states (the figure corresponds to the Néel and anti-Néel states) and evolved by applying a homogeneous quantum circuit, Trotterizing the integrable XXZ continuous-time dynamics ($\tau$ is the duration of the Trotter step). A lightcone spreads ballistically from the junction, inside of which non-trivial dynamics takes place. We study the profiles of local observables arising at late times. The plot shows the example of the staggered spin-$z$ magnetization and is obtained by solving the GHD equations, cf. Sec. 4.
Integrable quantum circuits providing a Trotterization of integrable Hamiltonians were introduced in Refs. [29, 30, 31]. In the past few years they have received significant attention due to their dynamical features [32, 33, 34, 35], their rich underlying mathematical structures [36, 37, 38, 39] and seminal work demonstrating experimental implementation in current quantum devices [40, 41, 42]. Prominently, Ref. [33] has shown the existence of Trotter transitions in integrable quantum circuits, manifesting in the correlations of the so-called generalized Gibbs ensemble (GGE) [26, 27, 28]. As GGEs describe the late-time physics of integrable systems, this result is interesting, as it implies that the information on the discretized structure of the dynamics is not washed out at coarse-grained time scales.
@@
The rest of this work is organized as follows. We begin in Sec. 2 introducing the XXZ quantum circuit and its Bethe ansatz solution. In Sec. 3 we discuss the bipartition protocol and derive the necessary technical ingredients for its analytic description. They include obtaining the GGEs corresponding to different initial states, generalizing the results of Ref. [33], and deriving the GHD equations for integrable quantum circuits. Our results are discussed in detail in Sec. 4, where we provide predictions for the system dynamics at different points in the parameter space, by numerically solving the GHD equations. Our conclusions are presented in Sec. 5, while the appendices contain the most technical aspects of our work.
2 The model
2.1 The integrable quantum circuit
@@
$V_{n,n+1}(\tau)=e^{-i\frac{\tau}{4}\left[\sigma_{n}^{x}\sigma_{n+1}^{x}+\sigma_{n}^{y}\sigma_{n+1}^{y}+\Delta(\sigma_{n}^{z}\sigma_{n+1}^{z}-\mathbb{1})\right]}\,,$ (4)
@@
$H=\frac{1}{4}\sum_{j=1}^{L}\left[\sigma^{x}_{j}\sigma^{x}_{j+1}+\sigma^{y}_{j}\sigma^{y}_{j+1}+\Delta(\sigma^{z}_{j}\sigma^{z}_{j+1}-\mathbb{1})\right]\,,$ (5)
@@
2.2 The Bethe ansatz solution
@@
The spectrum of $U(\tau)$ is organized into sectors labeled by the number $M$ of stable quasiparticles. The latter are magnonic excitations and, accordingly, $M$ is the quantum number associated with the conserved operator $\hat{M}=\sum_{j}(\mathbb{1}-\sigma_{j}^{z})/2$. In a given $M$-sector, the eigenstates are parametrized by sets of complex numbers $\{p_{j}\}_{j=1}^{M}$, satisfying the so-called Bethe equations [36]
@@ 2.2.1 The TBA in the gapped regime
$a_{nm}(\lambda)=$ $(1-\delta_{nm})a_{|n-m|}(\lambda)+2a_{|n-m|+2}(\lambda)+\ldots+2a_{n+m-2}(\lambda)+a_{n+m}(\lambda)\,,$ (14)
@@
$J_{2j}$ $={\mathcal{N}_{x}}^{-1}\left[2(\sigma_{2j-1}^{+}\sigma_{2j}^{-}-\sigma_{2j-1}^{-}\sigma_{2j}^{+})+i\sinh(ix)(\sigma_{2j-1}^{z}-\sigma_{2j}^{z})\right]\,,$ (22)
@@ 2.2.2 The TBA in the gapless regime
$\mathrm{sign}(q_{m})\left[\rho_{m}^{t}(\lambda)\right]=$ $a_{m}^{(ix/2)}(\lambda)-\sum_{n=1}^{N_{b}}\left(a_{mn}\ast\rho_{m}\right)(\lambda)\,,$ (32)
@@
$a_{jk}(\lambda)$ $=(1-\delta_{n_{j}n_{k}})a_{|n_{j}-n_{k}|}^{\upsilon_{j}\upsilon_{k}}(\lambda)+2a_{|n_{j}-n_{k}|+2}^{\upsilon_{j}\upsilon_{k}}(\lambda)+\ldots+2a_{n_{j}+n_{k}-2}^{\upsilon_{j}\upsilon_{k}}(\lambda)+a_{n_{j}+n_{k}}^{\upsilon_{j}\upsilon_{k}}(\lambda)\,.$ (35)
@@
$\mathrm{sign}(q_{n})\left[\rho_{n}^{t}(\lambda)v^{\rm eff}_{n}(\lambda)\right]=$ $\tfrac{1}{2\pi}\mathrm{sign}(q_{n})E^{\prime}_{n}(\lambda)-\sum_{m=1}^{N_{b}}\left(a_{nm}\ast(\rho_{m}v^{\rm eff}_{m})\right)(\lambda)\,,$ (36)
@@
$a^{(m)\,{\rm dr}}_{j}(\lambda)$ $=a_{p}(\lambda-(-1)^{m}\tfrac{ix}{2})-\sum_{q=1}^{N_{s}}\int{{\rm d}\mu}\,a_{jq}(\lambda-\mu)\theta_{q}(\mu)\sigma_{q}a^{(m)\,{\rm dr}}_{q}(\mu),$ (40)
@@
3 The bipartition protocol
3.1 The initial states
@@
Alternatively, one could consider preparing two halves of the system in non-equilibrium states with a simple structure [49]. As discussed in Sec. 1, each half of the system far enough from the junction is then expected to relax to a stationary GGE which depends on the initial state. Accordingly, at late time the non-trivial dynamics again takes place only in a region increasing linearly in time and centered around the junction.
@@ 3.1.1 Gapped phase
3.1.2 Gapless phase
As stated previously, the GGE description in the gapless phase depends in a non-trivial way on the specific value of $\gamma$. Following Section 2.2, we shall limit ourselves to the simplest case where $\gamma$ takes the form (28). The GGE is then characterized by a family of functions $\{\eta_{j}\}_{j=1,\ldots p+1}$, which are obtained from a set of auxiliary functions $\{Y_{j}\}_{j=1,\ldots p}$, $K$, $\mathfrak{b}$ whose construction is detailed in Appendix B. For concreteness, we give below the explicit rapidity distributions for $p=2,3$. For $p=2$ and the Néel state, we have
@@ 3.2 GHD equations
As we will discuss in the next subsection, the form of the scattering phase depends on the phase of the model and it can be read off by comparing  (55) with Eqs. (18) and (36), respectively.
3.3 The quasi-momentum and the quasi-energy
@@
3.3.1 Gapped phase
As discussed in Sec. 2, the eigenvalues of $U(\tau)$ depend on the solutions $\{p_{j}\}$ to the Bethe equation (8). Denoting by $u(\{p_{j}\})$ the eigenvalue of $U(\tau)$ corresponding to a set $\{p_{j}\}_{j=1}^{M}$, we have
@@ 3.3.1 Gapped phase
$E_{n}(\lambda)$ $=\!\sum_{j=-\tfrac{n-1}{2}}^{\tfrac{n-1}{2}}\varepsilon(i(\lambda+j\eta))=-\frac{i}{\tau}\!\log\quantity[\frac{\sin(\lambda-\tfrac{x}{2}+in\tfrac{\eta}{2})\sin(\lambda+\tfrac{x}{2}-in\tfrac{\eta}{2})}{\sin(\lambda-\tfrac{x}{2}-in\tfrac{\eta}{2})\sin(\lambda\!+\!\tfrac{x}{2}+in\tfrac{\eta}{2})}].$ (61)
@@
$E^{\prime}_{n}(\lambda)\!$ $=\!\frac{-\sinh({n\eta})/\tau}{\cosh^{2}(\tfrac{n\eta}{2})\sin^{2}(\lambda-\tfrac{x}{2})+\sinh^{2}(\tfrac{n\eta}{2})\cos^{2}(\lambda-\tfrac{x}{2})}-(x\to-x)\,.$ (62)
@@ 4 Results
where $\rho^{L}_{n}(\lambda)$, $\rho^{R}_{n}(\lambda)$ are the GGE distribution functions associated with the left and right initial states. For the product states considered in this work, these functions were computed in Sec. 3. The above boundary conditions encode the fact that, away from the junction point, the system will locally equilibrate to a homogeneous GGE, determined by the initial states.
Solving (71) with the boundary conditions (72) can be done with standard iterative methods, see e.g. Ref. [95]. In this work, we use a different numerical approach which was recently introduced in Ref. [96] and briefly explained in Appendix A. In the rest of this section we present the final results of our numerical computations. Note that the solution to (71) gives us a family of rapidity distribution functions $\{\rho_{n}(\zeta,\lambda)\}_{\zeta}$. We use the latter to compute the space-time profiles of local observables, using the formulas presented in Sec. 3.
4.1 The non-interacting points
@@
4.2 Observable profiles at the interacting points
@@ 5 Outlook
Acknowledgements
We acknowledge helpful discussions with Bruno Bertini and especially Benjamin Doyon for early collaboration on this project.
Funding information
FH acknowledges funding from the faculty of Natural, Mathematical & Engineering Sciences at King’s College London. The CREATE cluster [113] was used for numerical simulations. This work was co-funded by the European Union (ERC, QUANTHEM, 101114881). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.
@@ Appendix A Numerical solution of the GHD equation
We solve the GHD equation numerically using a recently introduced fixed point equation [96]. Given the states on the left (--) and the right side ($+$) either in terms of their quasi-particle densities $\rho_{\pm;i}(\lambda)$ or equivalently as occupation functions $\theta_{\pm;i}(\lambda)$, we can compute
@@
Then define:
$K_{0;i}(x,\lambda)$ $={P^{\prime}}^{\mathrm{dr}}_{\operatorname{sgn}(x),i}(\lambda)x$ (A.2)
$\hat{\theta}_{0;i}(k,p)$ $=\theta_{\operatorname{sgn}(k);i}(\lambda)$ (A.3)
$\hat{N}_{0;i}(k,p)$ $=\theta_{\operatorname{sgn}(k);i}(\lambda)k$ (A.4)
@@
$=P^{\prime}_{i}(\lambda)x-E_{i}^{\prime}(\lambda)t+\hat{\mathbf{T}}[\theta_{\operatorname{sgn}(K_{j}(t,x,\mu));j}(\mu)K_{j}(t,x,\mu)].$ (A.5)
@@
$\theta_{i}(t,x,\lambda)=\hat{\theta}_{i}(K_{i}(t,x,\lambda),\lambda)=\theta_{\operatorname{sgn}(K_{i}(t,x,\lambda));i}(\lambda).$ (A.6)
We solve Eq. (A.5) by iteration with initial guess $P^{\prime}_{i}(\lambda)x-E^{\prime}_{i}(\lambda)t$, which converges in the gappless phase, where only finitely many strings are present.
@@ A.1 Upgraded algorithm
As it turns out for the Néel and Dimer state in the gapped phase this iterative scheme does not converge. The same holds for the dressing equation (which can be viewed as a linearized version of (A.5)). The iteration scheme does not converge because $n_{i}(\lambda)$ does not decay sufficiently fast for large strings $i\to\infty$: Consider the first term of the iteration procedure to compute $f^{\mathrm{dr}}_{i}(\lambda)$
$f^{\mathrm{dr}}_{i}(\lambda)$ $=f_{i}(\lambda)+\hat{\mathbf{T}}\theta_{j}(\mu)f_{j}(\mu)+\ldots$ (A.7)
@@
$f_{i}^{\mathrm{dr}}(\lambda)$ $=f_{i}(\lambda)+\hat{\mathbf{T}}\theta_{j}(\mu)f_{j}^{\mathrm{dr}}(\mu).$ (A.8)
We can rewrite this as follows: Choose any functions $g_{i}(\lambda)$ and $m_{i}(\lambda)$ and define
$b_{i}(\lambda)=g_{i}(\lambda)-\hat{\mathbf{T}}m_{j}(\mu)g_{j}(\mu).$ (A.9)
Now write $f^{\mathrm{dr}}_{i}(\lambda)=w_{i}(\lambda)g_{i}(\lambda)$ and observe:
$w_{i}(\lambda)g_{i}(\lambda)$ $=f_{i}(\lambda)+\hat{\mathbf{T}}\theta_{j}(\mu)w_{j}(\lambda)g_{j}(\lambda)$ (A.10)
$w_{i}(\lambda)g_{i}(\lambda)$ $=w_{i}(\lambda)b_{i}(\lambda)+w_{i}(\lambda)\hat{\mathbf{T}}m_{j}(\mu)g_{j}(\mu).$ (A.11)
@@
$w_{i}(\lambda)$ $=\frac{1}{b_{i}(\lambda)}\left(f_{i}(\lambda)+\hat{\mathbf{T}}\theta_{j}(\mu)g_{j}(\lambda)w_{j}(\lambda)-w_{i}(\lambda)\hat{\mathbf{T}}m_{j}(\mu)g_{j}(\mu)\right).$ (A.12)
This is a new fixed point equation for $w_{i}(\lambda)$. It is motivated by the equation for the effective velocity, which is obtained in the special case $f_{i}(\lambda)=E^{\prime}_{i}(\lambda),b_{i}(\lambda)={P^{\prime}}_{i}(\lambda)$ and $m_{i}(\lambda)=n_{i}(\lambda)$. Intuitively speaking we need that $\zeta_{i}(\lambda)$ decays fast for large $i$, but at the same time $b_{i}(\lambda)$ should decay slowly, such that (A.12) converges.
For the gapped phase we used $g_{i}(\lambda)=1/2^{i}$ and $m_{i}(\lambda)=1$, which happened to make (A.12) convergent with initial guess $f_{i}(\lambda)/b_{i}(\lambda)$.
@@
$\tilde{K}_{i}(t,x,\lambda)$ $=\tfrac{1}{b_{i}(\lambda)}\Big{(}P^{\prime}_{i}(\lambda)x-E_{i}^{\prime}(\lambda)t-\tilde{K}_{i}(t,x,\lambda)\hat{\mathbf{T}}m_{j}(\mu)g_{j}(\mu)+\hat{\mathbf{T}}[\hat{N}_{0;j}(g_{j}(\mu)\tilde{K}_{j}(t,x,\mu),\mu)]\Big{)},$ (A.13)
@@
$\tilde{K}_{i}(t,x,\lambda)$ $=\tfrac{1}{b_{i}(\lambda)}\Big{(}P^{\prime}_{i}(\lambda)x-E_{i}^{\prime}(\lambda)t-\tilde{K}_{i}(t,x,\lambda)\hat{\mathbf{T}}m_{j}(\mu)g_{j}(\mu)$
$+$ $\hat{\mathbf{T}}[\theta_{\operatorname{sgn}(\tilde{K}_{j}(t,x,\mu));j}(\mu)g_{j}(\mu)\tilde{K}_{j}(t,x,\mu)]\Big{)}.$ (A.14)